1)     In a frequency distribution of 100 families given below the number of families corresponding to expenditure groups 20-40 and 60-80 are missing from the table. However, the median is known to be 50.

 Expenditure group

0-20

20-40

40-60

60-80

80-100

No. of families

14

f2

27

f4

15

a)         Find the missing frequencies.

b)        Find mean, mode, variance, standard deviation, range, mean deviation for mean and median


1
Expert's answer
2021-01-29T14:06:37-0500

a) Let "f_2=" the number of families in class "20-40," "f_4=" the number of families in class

"60-80." Then


"14+f_2+27+f_4+15=100"

"f_2+f_4=44"

Estimated Median ="L+\dfrac{n/2-B}{G}\times w"  

where:

L is the lower class boundary of the group containing the median

n is the total number of values

B is the cumulative frequency of the groups before the median group

G is the frequency of the median group

w is the group width

Given "n=100, w=20, Median=M=50"

Then "L=20" or "L=40"


Take "L=20"


"20+\dfrac{100/2-14}{f_2}\times 20=50"

"f_2=24"

But "20+24=44<100\/2." We have a contradiction.


Take "L=40"


"40+\dfrac{100/2-(14+f_2)}{27}\times 20=50"

"36-f_2=13.5"

"f_2=22.5"

Take "f_2=23. f_4=21"


b)


"10(14)+30(23)+50(27)+70(21)+90(15)=5000"


"mean=\dfrac{5000}{100}"

"mean=\bar{x}=50"

Find Mode Class

Maximum frequency is 27.The mode class is 40-60.

"L=40"

"f_3=27"

"f_2=23"

"f_4=21"

"w=20"


"mode=L+\dfrac{f_3-f_2}{2f_3-f_2-f_4}\times w"

"=40+\dfrac{27-23}{2(27)-23-21}\times 20=48"

"10^2(14)+30^2(23)+50^2(27)+70^2(21)+90^2(15)="

"=314000"

"314000-\dfrac{(5000)^2}{100}=64000"

"Var(X)=s^2=\dfrac{64000}{100-1}=646.4646"

"s=\sqrt{s^2}=\sqrt{\dfrac{64000}{99}}=25.4257"

Mean deviation of Mean


"\delta\bar{x}=\dfrac{\sum_i|x_i-\bar{x}|f_i}{100}"

"\sum_i|x_i-\bar{x}|f_i""=40(14)+20(23)+0(27)+20(21)+40(15)=2040"

"\delta\bar{x}=\dfrac{\sum_i|x_i-\bar{x}|f_i}{100}=\dfrac{2040}{100}=20.4"

Mean deviation of Median


"\delta M=\dfrac{\sum_i|x_i-M|f_i}{100}"

Since "\bar{x}=M"

"\delta M=\delta\bar{x}=20.4"


"Range=90-10=80"